A machine produces cylindrical metal pieces. A sample of pieces is taken, and the diameter are found to be 1.01, 0.97, 1.03, 1.04, 0.99, 0.98, 0.99, 1.01 and 1.03 cm. Find the t-value for a 98% confidence interval for the diameters of the cylindrical metal pieces. Assume that the diameters are approximately normally distributed.


1
Expert's answer
2022-04-12T16:44:11-0400

Since the population value of σ is unknown, we want to use the confidence interval for σ unknown. Thus the 1 − α confidence interval is

"\bar{x}-t_{\alpha/2}\cfrac{s}{\sqrt{n}}<\mu<\bar{x}+t_{\alpha/2}\cfrac{s}{\sqrt{n}}"

"\bar{x}-t_{\alpha/2}\cfrac{s}{\sqrt{n}}<\mu<\bar{x}+t_{\alpha/2}\cfrac{s}{\sqrt{n}}""\bar{x}-t_{\alpha/2}\cfrac{s}{\sqrt{n}}<\mu<\bar{x}+t_{\alpha/2}\cfrac{s}{\sqrt{n}}"

where "t_{\alpha/2}" is the value from the t-distribution table for

"P(T>t_{\alpha/2})=\alpha/2."

For a 98% confidence interval we have α = 0.02 and α/2 = 0.01.

Using the table for df = n - 1 = 9 - 1 = 8 (degrees of freedom, n = 9 - the sample size):

"t_{0.01;8}=2.896."

The population sample mean:

"\bar{x}=(1.01+0.97+1.03+1.04+0.99+0.98+\\+0.99+1.01+1.03)/9=1.0056."

The population variance:

"s^2=\cfrac{\sum(x_i-\bar x)^2}{n-1}=\\
=((1.01-1.0056)^2+(0.97-1.0056)^2+\\
+(1.03-1.0056)^2+(1.04-1.0056)^2+\\
+(0.99-1.0056)^2+(0.98-1.0056)^2+\\
+(0.99-1.0056)^2+(1.01-1.0056)^2+\\
+(1.03-1.0056)^2)/8=0.00060278."

The population standard deviation:

"s=\sqrt{0,00060278}=0.0246."

"s=\sqrt{0,00060278}=0.0246."

Thus the 98% CI is

"1.0056-2.896\cdot\cfrac{0.0246}{\sqrt{9}}<\mu<1.0056+2.896\cdot\cfrac{0.0246}{\sqrt{9}}\\
0.9819<\mu<1.0293."


"1.0056-2.896\cdot\cfrac{0.0246}{\sqrt{9}}<\mu<1.0056+2.896\cdot\cfrac{0.0246}{\sqrt{9}}\\
0.9819<\mu<1.0293.""1.0056-2.896\cdot\cfrac{0.0246}{\sqrt{9}}<\mu<1.0056+2.896\cdot\cfrac{0.0246}{\sqrt{9}}\\
0.9819<\mu<1.0293."